OpenStudy (lisa123):

A sheet of aluminum foil (25 in. x 25 in.) weighs 5 oz. and is 0.20 in. thick. What is the density (in g/cm3) of the sheet of aluminum foil?

OpenStudy (lisa123):

Please help @dan815 and @zepdrix

OpenStudy (mathstudent55):

You need the density in g/cm^3, but you are given the dimensions in in., and the weight or mass in oz. You will need to convert units.

OpenStudy (mathstudent55):

First, since the density you want is in g/cm^3, that means you need the weight or mass and the volume. You are given 5 oz. Do you know how to calculate the volume from the given dimensions, 25 in. x 25 in. x 0.20 in.?

OpenStudy (lisa123):

okay I think I got it! My final answer is 0.069 g/cm^3

OpenStudy (mathstudent55):

\(\large \dfrac{5~oz}{25~in. \times 25~in. \times 0.20~in.} = 0.04 \dfrac{oz}{in.^3}\) \(\large 0.04 \dfrac{oz}{in.^3} \times \dfrac{in.^3}{(2.54)^3 cm^3} \times \dfrac{1~lb}{16~oz} \times \dfrac{453.59237~g}{1~lb} = 0.069 \dfrac{g}{cm^3}\)

OpenStudy (mathstudent55):

I get the same as you. The density of aluminum is 2.7 g/cm^3, so I don't really understand this problem. Are you sure the numbers you wrote above are correct?

OpenStudy (mathstudent55):

0.20 in. is 5.08 mm. That is more than 0.5 cm. That is an impossible thickness for aluminum foil. Aluminum foil may be 0.2 mm thick, not 0.2 in. thick. The problem is incorrect as it is posted.